# Combined Pre-Supernova Alert System with KamLAND and Super-Kamiokande

### v2

@@ 4 The Super-Kamiokande experiment
The SK experiment is a water Cherenkov detector located in the Kamioka mine, the same location as KamLAND-Zen, overburdened with 1,000 m of rock (equivalent to 2,700 meters of water) to reduce cosmic-ray muon backgrounds. SK is a multi-purpose detector, which has been operating since 1996 and focuses on nucleon decays and neutrino properties such as neutrino oscillations by observing atmospheric, solar, and accelerator neutrinos. Furthermore, SK is a neutrino telescope, capable of observing neutrinos emitted by stars and the universe as a whole.
The SK detector is composed of a cylindrical stainless steel tank with 39.3 m diameter and 41.4 m height (Fukuda et al., 2003). It contains 50 kton of water doped with Gd since 2020 (Abe et al., 2022a). The detector is divided into two regions: the inner detector (ID) and the outer detector (OD). The ID is responsible for the event detection, with over 11,000 20-inch PMTs and it has a volume of 32 kton, although the usual FV used in SK analyses is 22.5 kton. The OD has a thickness of about 2 m and it is composed of 1,885 8-inch PMTs, facing the outside of the detector to reduce entering cosmic-ray muon induced backgrounds.
In 2020, Gd sulfate octahydrate was loaded to the water in the detector, starting the SK-Gd phase. The loading of Gd improves SK’s sensitivity to low-energy , expanding the physics goals of the experiment. SK has now the potential to reveal neutrinos from the Diffuse Supernova Neutrino Background (DSNB) (Abe et al., 2021; Harada et al., 2023) and pre-SN stars, which are yet to be observed. In 2022, an additional Gd loading into SK was completed, achieving higher concentrations of Gd in the water.
The detection of low-energy from pre-SN stars in SK is through IBD interaction, same as in KamLAND. However, positrons generated from IBD produce Cherenkov light instead of scintillation light. And -rays from neutron capture are detected mainly by Compton scattering electrons, producing Cherenkov light. In SK-Gd, the majority of thermal neutron capture will be on Gd because of the overwhelmingly large neutron capture cross section: while the capture cross section in hydrogen is only 0.3 barns, the average effective cross section in Gd is 49,000 barns. The largest contributions for the neutron capture come from the isotopes 157Gd and 155Gd (Abe et al., 2022a). The -ray cascade produced by the thermal neutron capture on Gd (nGd) has a higher energy than the single -ray produced by capture on hydrogen, therefore many more photons are detected. The energy of the cascade is around 8 MeV. In the first phase of SK-Gd (July 2020-March 2022), which corresponded to a concentration of Gd by mass, approximately of neutron captures were on Gd. For the current phase with Gd by mass (since July 2022), the neutron capture efficiency is approximately .
@@ 4.1 Event selection in SK
The pre-SN alert system receives data from WIT right after the organizer processes sorts the data in time. The system runs the event selection in real time, which is based on the coincidence distance () and coincidence time () of IBD pairs and two Boosted Decision Tree (BDT) methods: one used as pre-selection for IBD pair identification (BDTonline) and another used as final selection based on angular distribution of hits, reconstructed energy and quality (BDToffline) (more details in (Machado et al., 2022)). For the current SK-Gd phase with Gd, BDTs were re-trained and cuts were optimized: BDT, cm, s, and BDT. Figure 3 shows the signal background separation of the current BDTonline used for pre-selection and Figure 4 shows the efficiency of applying the selection criteria to 107 IBD pairs. The irreducible background rate is approximately 0.5 event/hour.
@@ 4.2 Background sources in SK
The major backgrounds for pre-SN neutrino search in SK are reactor neutrinos and accidental coincidences. Other background sources include geoneutrinos, radioactive contaminants, and cosmic-ray muon induced spallations. The backgrounds from geoneutrinos and accidental coincidences are much similar to what has been described for KamLAND in Section 3.2. Reactor neutrino background will be discussed in Section 5.2. Radioactive contaminants that came along with the Gd loading are also a background source: 235U chain isotopes can emit , contributing to the backgrounds from 18O(,n)21Ne∗ and 17O(,n)20Ne∗ processes. Moreover, the spontaneous fission of 238U can emit neutrons that mimic delayed signals or even IBD candidates. Cosmic-ray muon induced spallations are expected to have a tiny contribution since its resulting background rate is low, and are effectively removed using BDTs.
@@ 5 Sensitivity to pre-SN neutrinos in KamLAND and SK
Previous sensitivity studies for detection of pre-SN neutrinos in KamLAND and SK are given in (Asakura et al., 2016; Simpson et al., 2019; Machado et al., 2022). In (Asakura et al., 2016), the sensitivity for KamLAND’s detection of pre-SN neutrinos assuming the model from (Odrzywolek & Heger, 2010) is presented. (Simpson et al., 2019) presents a preliminary overall sensitivity for SK doped with Gd by mass. Later in (Machado et al., 2022), an improved sensitivity to pre-SN neutrinos in SK is shown, using data from the first phase of SK-Gd (with Gd) to predict realistic backgrounds and new event selection methods.
@@ 5.1 Analysis strategies
The general analysis strategy for each experiment is as follows. A rapid increase in the candidate event rate is sought without explicit reference to any of the pre-SN neutrino models. In each experiment, the background rate is measured over a relatively long period (30 days or more) using recent data, in order to reduce the effects of random fluctuations in the data. An analysis window of a few hours is used to measure the observed event rate for the purpose of searching for signal events. The detection significance is calculated by comparing the observed event rate to the expected background rate. In other words, a test of significance is performed, with the null hypothesis being that the observed event rate is consistent with the expected background rate within the sampling error.
KamLAND and SK are in rather different experimental conditions, such as target mass, detection energy threshold, background rates, and duty cycle, etc. These factors affect the choices of background time window and analysis time window. SK chooses a 30-day background time window, while KamLAND’s choice is a longer 90-day background time window due to its lower background rate. The analysis time windows are chosen based on the principle of achieving the longest warning time. As a result, KamLAND has chosen an optimal time window of 24 hours. For SK, the time window was optimized to 12 hours, maximizing the warning time for Betelgeuse-like models and reducing the impact that potential interruption in DAQ and calibration work in the detector has in the pre-SN alert system.
To evaluate the significance, the following steps are applied. Both experiments are considered as Poisson counting experiments. Let be the number of candidates observed in the analysis time window at time . Since only an excess of the event rate appears to be signal-like, the -value at is defined as
where is the measured number of background in the background time window and is normalized to the analysis time window. The significance is given by
where is the cumulative standard normal distribution function.
@@ 5.2 Background assumptions
The background rates in KamLAND and SK can be largely affected by the nuclear reactors near the Kamioka mine. Reactor are mainly generated through the beta decays of the fission products of 235U, 238U, 239Pu and 241Pu in nuclear reactors. The Great East Japan Earthquake of 2011 caused numerous reactors in Japan to be shut down, and they have been gradually restarting since 2015. A constant monitoring of the situation in Japan’s nuclear power plants is being conducted, and backgrounds are being updated. Nonetheless, it is difficult to predict when and which reactors will start operation in the future. Therefore, we assume different reactor background conditions according to three scenarios: low, medium and high reactor activities. The low reactor activity scenario assumes that all reactors in Japan are not in operation. For the other two scenarios, reactors near the Kamioka mine (Mihama 3 unit, Ohi 3, 4 units and Takahama 1, 2, 3, 4 units) are assumed to be operating with a 100% load factor for medium activity and with 200% load factor for high activity. We note that medium reactor activity is close to the situation as of the winter of 2023-2024.
Figure 5 shows the expected reactor fluxes at the Kamioka mine considering these three reactor scenarios, as well as the expected geoneutrino fluxes. The fluxes are calculated based on the spectra per fission from (Huber, 2011; Mueller et al., 2011; Vogel et al., 1981), with relative fission yields (235U,238U,239Pu,241Pu) assumed be (0.570,0.078,0.295,0.057) (Eguchi et al., 2003). And the values of neutrino oscillation parameters are eV2, , and . The thermal power of each reactor is estimated from publicly available data on reactor electric power. In the low reactor activity scenario, where all reactors in Japan are assumed to be off, the only contribution we consider comes from reactors in Korea. In addition to the reactor neutrino fluxes, we also plot the geoneutrino flux at the Kamioka mine for comparison. The geoneutrino flux is calculated based on the model in (Enomoto et al., 2007). For energy below 2.2 MeV, the geoneutrino flux is comparable to the reactor neutrino flux under the high reactor activity assumption. It decreases rapidly at 2.2 MeV, and is roughly the same as the reactor neutrino flux assuming low reactor activity between 2.2 MeV and 3 MeV. Above 3 MeV, the geoneutrino flux becomes negligible.
@@ 5.3 False alarm rate
It is a common practice to quantify the false positive rate of a statistical test using the -value of the test. However, in this search, we find it misleading to report the results using the -value or the significance of a single test. The online search of pre-SN neutrinos is performed continuously, while the time when the pre-SN neutrino signal appears is not known in advance. An appropriate way to estimate the -value is to figure out the probability, assuming background-only, of finding a signal at any time that is at least as extreme as the one observed. Such a -value can be substantially higher than the -value of a single test. This is the so-called “look-elsewhere effect” (Lyons, 2008).
To resolve this issue, we introduce the quantity “false alarm rate” to report the result of the search. The false alarm rate is the expected frequency that a false alarm happens in units of per century. A false alarm is caused by a significant signal due to background fluctuations. The false alarm rate can be derived from toy Monte-Carlo simulations, assuming a background-only scenario. In practice, we generate a time series of Poisson random events with the expected value being the background rate. The statistical test described in the previous subsection is performed. Then we evaluate the frequency with which the random events cause a significant signal. Thereby, a false alarm rate is found. Considering the frequency of CCSN in the galaxy is approximately once every few decades (Adams et al., 2013), we set a as the ultimate pre-SN alert criteria.
@@ 5.4 Sensitivity to pre-SN neutrinos at KamLAND
We evaluate the expected numbers of signal events in KamLAND for the two pre-SN neutrino models with neutrino oscillation effects, assuming a Betelgeuse-like pre-SN star. Figure 6 (a) shows the expected number of signal events, integrated over a 24-hour time window, as a function of time to CCSN. Figure 6 (b) shows the integrated number of signals for the last 24 hours before core collapse at different distances. The expected background counts integrated over 24 hours is for low, medium, and high reactor activities, respectively. The background rate in KamLAND is sufficiently low, such that a few candidate events can cause a significant signal.
Figure 8 presents the time evolution of the expected detection significance assuming medium reactor activity. The significance corresponding to are also plotted as horizontal dotted-dashed lines. The remaining time to core collapse, when KamLAND observes an excess of pre-SN neutrino candidates as extreme as , is defined as the warning time. Note that the latency due to data processing is not taken into account when calculating the warning time. For a Betelgeuse-like pre-SN star, KamLAND is capable of issuing a pre-SN alert 6.5 hours before the CCSN, assuming the Odrzywolek model and normal mass ordering. In the case of inverted mass ordering, the warning time is largely shortened as the pre-SN neutrino fluxes become lower. The worst case is, with the Odrzywolek model and inverted mass ordering, the expected detection significance cannot reach the alert criteria. The above discussions are based on the medium reactor activity scenario. Table 2 summarizes the warning time for all three reactor activity assumptions. If the reactor activity assumption shifts from medium to high, for the normal ordering cases, the warning time will be shortened by roughly 1 hour. For the inverted ordering cases, with the high reactor activity assumption, KamLAND is unable to issue an alert with a , noted as “N/A (Not Applicable)” in Table 2.
Figure 8 pictures the warning time as a function of distance. The lines are estimations assuming medium reactor activity. The upper edges of the bands are for low reactor activity, and the lower edges are for high reactor activity. These results indicate that, for nearby pre-SN candidates, KamLAND can send alerts tens of hours before the explosion. And for the medium reactor activity case, KamLAND is sensitive to pre-SN candidates within an optimistic distance of 280 pc away from Earth.
@@ 5.5 Sensitivity to pre-SN neutrinos at SK
The fluxes of pre-SN neutrinos are taken from the two pre-SN models, with neutrino oscillation effects assuming normal and inverted mass orderings. The expected signal rates are aggregated over a 12-hour time window, resulting in the expected number of signal events as a function of time, as pictured in Figure 9 (a). Figure 9 (b) presents the number of signals integrated over the last 12 hours for different distances. The expected background counts for low, medium, and high reactor activities in SK are 4.6, 6.2 and 8.1 events. Although the background rate is much higher than that in KamLAND, the large target volume allows SK to collect signal events an order of magnitude more than KamLAND does. Thus the statistical significance in SK can increase rapidly when approaching core collapse.
Based on the estimations of signal and background, we assessed the sensitivity of pre-SN neutrino detection in SK-Gd with Gd loaded. Figure 10 presents the time evolution of the expected detection significance in SK-Gd, assuming medium reactor activity. The results show that SK-Gd is capable of providing an early warning before the CCSN, at most 10.9 hours assuming the Patton model and normal ordering, and at least 2.1 hours for the Odrzywolek model and inverted ordering. The warning time for all of the simulated scenarios are summarized in Table 2. If the reactor neutrino fluxes around the Kamioka mine are doubled, the warning time can be shortened by 0.2-1.1 hours.
We plot the warning time as a function of distance in Figure 11. The upper (lower) edges of the bands are for low (high) reactor activity, and the lines in between are for medium reactor activity. Under neutrino flux assumptions of Patton model and normal ordering, the SK alert can cover 15 M⊙ stars with a distance of 500 pc from Earth, for the medium reactor activity case.

### v3

@@ 4 The Super-Kamiokande experiment
The SK experiment is a water Cherenkov detector located in the same Kamioka mine as KamLAND. SK is a multi-purpose detector, which has been operating since 1996 and focuses on nucleon decays (Takenaka et al., 2020) and neutrino properties such as neutrino oscillations by observing atmospheric (Wester et al., 2024), solar (Abe et al., 2024), and accelerator neutrinos (Abe et al., 2023a). Furthermore, SK is a neutrino telescope, capable of observing neutrinos from astronomical sources (Abe et al., 2021a; Orii et al., 2021; Mori et al., 2022).
The SK detector is composed of a cylindrical stainless steel tank with 39.3 m diameter and 41.4 m height (Fukuda et al., 2003). The detector is divided into two regions: the inner detector (ID) and the outer detector (OD). The ID is responsible for the event detection, with over 11,000 20-inch PMTs and it has a volume of 32 kton, although the usual FV used in SK analyses is 22.5 kton. The OD has a thickness of about 2 m and it is composed of 1,885 8-inch PMTs, facing the outside of the detector to reduce entering cosmic-ray muon induced backgrounds.
In 2020, Gd sulfate octahydrate was dissolved to the water in the detector, starting the SK-Gd phase. The loading of Gd improves SK’s sensitivity to low-energy , expanding the physics goals of the experiment. SK has now the potential to reveal neutrinos from the Diffuse Supernova Neutrino Background (DSNB) (Abe et al., 2021b; Harada et al., 2023) and pre-SN stars, which are yet to be observed. In 2022, an additional Gd loading into SK was completed, achieving higher concentrations of Gd in the water (Abe et al., 2022a).
Low-energy from pre-SN stars are detected in SK via IBD, similarly to KamLAND. However, positrons generated from IBD produce Cherenkov radiation instead of scintillation light, and -rays from neutron capture are detected mainly by Compton scattering electrons, producing Cherenkov radiation. In SK-Gd, the majority of thermal neutron captures occur on Gd due to its significantly higher neutron capture cross-section. Specifically, while hydrogen has a capture cross-section of only 0.3 barns, Gd’s effective cross-section averages 49,000 barns. The largest contributions for the neutron capture come from the isotopes 157Gd and 155Gd (Abe et al., 2022a). The resulting -ray cascade from neutron captures on Gd (nGd) releases more energy—approximately 8 MeV—compared to captures on hydrogen, leading to a greater photon yield. In the first phase of SK-Gd (July 2020-March 2022), which corresponded to a concentration of Gd by mass, approximately of neutron captures were on Gd. For the current phase with Gd by mass (since July 2022), the neutron capture efficiency is approximately .
@@ 4.1 Event selection in SK
The pre-SN alert system receives data from WIT right after the organizer processes sort the data in time. The system runs the event selection in real time, which is based on the coincidence distance () and coincidence time () of IBD pairs and two Boosted Decision Tree (BDT) methods: one used as pre-selection for IBD pair identification (BDTonline) and another used as final selection based on angular distribution of hits, reconstructed energy and quality (BDToffline) (more details in (Machado et al., 2022)). For the current SK-Gd phase with Gd, BDTs were re-trained and cuts were optimized: BDT, cm, s, and BDT. Figure 3 shows the signal background separation of the current BDTonline used for pre-selection and Figure 4 shows the efficiency of applying the selection criteria to 107 IBD pairs. The irreducible background rate is approximately 0.5 event/hour.
@@ 4.2 Background sources in SK
The major backgrounds for the pre-SN neutrino search in SK are reactor neutrinos and accidental coincidences. Other background sources include geoneutrinos, radioactive contaminants, and cosmic-ray muon induced spallations. The backgrounds from geoneutrinos and accidental coincidences are similar to what has been described for KamLAND in Section 3.2. Reactor neutrino background will be discussed in Section 5.2. Radioactive contaminants that came along with the Gd loading are also a background source: 235U chain isotopes can emit , contributing to the backgrounds from 18O(,n)21Ne∗ and 17O(,n)20Ne∗ processes. Moreover, the spontaneous fission of 238U can emit neutrons that mimic delayed signals or even IBD candidates. Cosmic-ray muon induced spallations are expected to have a tiny contribution since its resulting background rate is low, and are effectively removed using BDTs.
@@ 5 Sensitivity to pre-SN neutrinos in KamLAND and SK
Previous sensitivity studies for detection of pre-SN neutrinos in KamLAND and SK are given in (Asakura et al., 2016; Simpson et al., 2019; Machado et al., 2022). In (Asakura et al., 2016), the sensitivity for KamLAND’s detection of pre-SN neutrinos assuming the model from (Odrzywolek & Heger, 2010) is presented. (Simpson et al., 2019) presents a preliminary overall sensitivity for SK doped with Gd by mass. In (Machado et al., 2022), an improved sensitivity to pre-SN neutrinos in SK is shown, using data from the first phase of SK-Gd (with Gd) to predict realistic backgrounds and new event selection methods.
@@ 5.1 Analysis strategies
The general analysis strategy for each experiment is as follows. A rapid increase in the candidate event rate is sought without explicit reference to any of the pre-SN neutrino models. In each experiment, the background rate is measured over a relatively long period (30 days or more) using recent data, in order to reduce the effects of random fluctuations in the data. A sliding analysis window of a few hours is used to measure the observed event rate for the purpose of searching for signal events. The detection significance is calculated by comparing the observed event rate to the expected background rate. In other words, a test of significance is performed, with the null hypothesis being that the observed event rate is consistent with the expected background rate within the sampling error.
Although KamLAND and SK are at nearly identical locations, they are in rather different experimental conditions, such as target mass, detection energy threshold, background rates, and duty cycle, etc. These factors affect the choices of background time window and analysis time window. SK chooses a 30-day background time window, while KamLAND’s choice is a longer 90-day background time window due to its lower background rate. The analysis time windows are chosen based on the principle of achieving the longest warning time. As a result, KamLAND has chosen an optimal time window of 24 hours. For SK, the time window was optimized to 12 hours, maximizing the warning time for Betelgeuse-like models and reducing the impact that potential interruptions in DAQ and calibration work in the detector may have in the pre-SN alert system.
In this analysis, both experiments are considered as Poisson counting experiments, and their Poisson likelihoods and are constructed. The subscripts SK and KL denote Super-Kamiokande and KamLAND, respectively. The Poisson likelihood for each experiment can be written as
where the subscript can be SK or KL. is the observed number of events within the sliding analysis time window. The term , being the expected number of events, is given by
where is the parameter for the number of signal and is the expected number of background. The test statistic based on the likelihood ratio is given by
where the numerator and denominator are the maximum likelihoods with and without imposing a background-only scenario (Cowan et al., 2011). The more the observation disagrees with the background-only hypothesis, the larger is. As the test statistic asymptotically approaches distribution, we consider as the detection significance for each experiment.
@@ 5.2 Background assumptions
The background rates in KamLAND and SK can be largely affected by the nuclear reactors near the Kamioka mine. Reactor are mainly generated through the beta decays of the fission products of 235U, 238U, 239Pu and 241Pu in nuclear reactors. The Great East Japan Earthquake of 2011 caused numerous reactors in Japan to be shut down, and they have been gradually restarting since 2015. A constant monitoring of the situation in Japan’s nuclear power plants is being conducted, and backgrounds are being updated. Nonetheless, it is difficult to predict when and which reactors will start operation in the future. Therefore, we assume different reactor background conditions according to three scenarios: low, medium and high reactor activities. The low reactor activity scenario assumes that all reactors in Japan are not in operation. For the medium reactor activity scenario, reactors near the Kamioka mine (Mihama 3 unit, Ohi 3, 4 units and Takahama 1, 2, 3, 4 units) are assumed to be operating with a 100% load factor. We note this is close to the situation as of the winter of 2023-2024. The high reactor activity scenario assumes the amount of reactor neutrinos is doubled compared to the medium reactor activity scenario.
Figure 5 shows the expected reactor fluxes at the Kamioka mine considering these three reactor scenarios, as well as the expected geoneutrino fluxes. The fluxes are calculated based on the spectra per fission from (Huber, 2011; Mueller et al., 2011; Vogel et al., 1981), with relative fission yields (235U,238U,239Pu,241Pu) assumed be (0.570,0.078,0.295,0.057) (Eguchi et al., 2003). The values of neutrino oscillation parameters are eV2, , and . The thermal power of each reactor is estimated from publicly available data on reactor electric power. In the low reactor activity scenario, where all reactors in Japan are assumed to be off, the only contribution we consider comes from reactors in Korea. In addition to the reactor neutrino fluxes, we also plot the geoneutrino flux at the Kamioka mine for comparison. The geoneutrino flux is calculated based on the model in (Enomoto et al., 2007). For energy below 2.2 MeV, the geoneutrino flux is comparable to the reactor neutrino flux under the high reactor activity assumption. It decreases rapidly at 2.2 MeV, and is roughly the same as the reactor neutrino flux assuming low reactor activity between 2.2 MeV and 3 MeV. Above 3 MeV, the geoneutrino flux becomes negligible.
@@ 5.3 False alarm rate
It is a common practice to quantify the false positive rate of a statistical test using the -value of the test. However, in this search, we find it misleading to report the results using the -value or the significance of a single test. The online search of pre-SN neutrinos is performed continuously, while the time when the pre-SN neutrino signal appears is not known in advance. An appropriate way to estimate the -value is to determine the probability, assuming background-only, of finding a signal at any time that is at least as extreme as the one observed. Such a -value can be substantially higher than the -value of a single test. This is the so-called “look-elsewhere effect” (Lyons, 2008).
To resolve this issue, we introduce the quantity “false alarm rate” to report the result of the search. The false alarm rate is the expected frequency that a false alarm may happen per century. A false alarm is caused by a significant signal due to background fluctuations. The false alarm rate can be derived from toy Monte-Carlo simulations, assuming a background-only scenario. In practice, we generate a time series of Poisson random events with the expected value being the background rate. The statistical test described in the previous subsection is performed. Then we evaluate the frequency with which the random events cause a significant signal. Thereby, a false alarm rate is found. Considering the frequency of CCSN in the galaxy is approximately once every few decades (Adams et al., 2013), we set a as the ultimate pre-SN alert criteria.
@@ 5.4 Sensitivity to pre-SN neutrinos at KamLAND
We evaluate the expected numbers of signal events in KamLAND for the two pre-SN neutrino models with neutrino oscillation effects, assuming a Betelgeuse-like pre-SN star. Figure 6 (a) shows the expected number of signal events, integrated over a sliding 24-hour time window, as a function of time to CCSN. Figure 6 (b) shows the integrated number of signals for the last 24 hours before core collapse at different distances. The expected background counts integrated over 24 hours are for low, medium, and high reactor activities, respectively. The background rate in KamLAND is sufficiently low, such that a few candidate events can cause a significant signal.
Figure 8 presents the time evolution of the expected detection significance assuming medium reactor activity. The significance corresponding to false alarm rates of 1, 10, and 100 per century are also plotted as horizontal dotted-dashed lines. The remaining time to core collapse, when KamLAND observes an excess of pre-SN neutrino candidates as extreme as , is defined as the warning time. Note that the latency due to data processing is not taken into account when calculating the warning time. For a Betelgeuse-like pre-SN star, KamLAND is capable of issuing a pre-SN alert 6.5 hours before the CCSN, assuming the Odrzywolek model and normal mass ordering. In the case of inverted mass ordering, the warning time is largely shortened as the pre-SN neutrino fluxes become lower. The worst case is, with the Odrzywolek model and inverted mass ordering, the expected detection significance cannot reach the alert criteria. The discussions above are based on the medium reactor activity scenario. Table 2 summarizes the warning time for all three reactor activity assumptions. If the reactor activity assumption shifts from medium to high, for the normal ordering cases, the warning time will be shortened by roughly 1 hour. For the inverted ordering cases, with the high reactor activity assumption, KamLAND is unable to issue an alert with a , noted as “N/A (Not Applicable)” in Table 2.
Figure 8 pictures the warning time as a function of distance. The lines are estimations assuming medium reactor activity. The upper edges of the bands are for low reactor activity, and the lower edges are for high reactor activity. These results indicate that, for nearby pre-SN candidates, KamLAND can send alerts tens of hours before the explosion. For the medium reactor activity case, KamLAND is sensitive to pre-SN candidates within an optimistic distance of 280 pc away from Earth.
@@ 5.5 Sensitivity to pre-SN neutrinos at SK
The fluxes of pre-SN neutrinos are taken from the two pre-SN models, with neutrino oscillation effects assuming normal and inverted mass orderings. The expected signal rates are aggregated over a sliding 12-hour time window, resulting in the expected number of signal events as a function of time, as pictured in Figure 9 (a). Figure 9 (b) presents the number of signals integrated over the last 12 hours for different distances. The expected background counts for low, medium, and high reactor activities in SK are 4.6, 6.2 and 8.1 events. Although the background rate is much higher than that in KamLAND, the large target volume allows SK to collect signal events an order of magnitude larger than KamLAND. Thus the statistical significance in SK can increase rapidly when approaching core collapse.
Based on the estimations of signal and background, we assessed the sensitivity of pre-SN neutrino detection in SK-Gd with Gd loading. Figure 10 presents the time evolution of the expected detection significance in SK-Gd, assuming medium reactor activity. The results show that SK-Gd is capable of providing an early warning before the CCSN, at most 10.9 hours assuming the Patton model and normal ordering, and at least 2.1 hours for the Odrzywolek model and inverted ordering. The warning time for all of the simulated scenarios are summarized in Table 2. If the reactor neutrino fluxes around the Kamioka mine are doubled, the warning time can be shortened by 0.2-1.1 hours.
We plot the warning time as a function of distance in Figure 11. The upper (lower) edges of the bands are for low (high) reactor activity, and the lines in between are for medium reactor activity. Under neutrino flux assumptions of the Patton model and normal ordering, the SK alert can cover 15 M⊙ stars with a distance of 500 pc from Earth, for the medium reactor activity case.